OpenStudy (anonymous):

What is the slope of the line that is perpendicular to the line whose equation is 2x + y = 4. A.) -2 B.) -1/2 C.) 1/2 D.) 2

OpenStudy (ajprincess):

2x + y = 4. can u pls tell me vat u get when u solve the above equation for y?

OpenStudy (anonymous):

What is true about the slopes of perpendicular lines?

OpenStudy (anonymous):

2?

OpenStudy (ajprincess):

nope:( well have u solved the equation for y?

OpenStudy (anonymous):

yea and i got 2

OpenStudy (ajprincess):

2? can u pls show me ur work?

OpenStudy (ajprincess):

@DvonCruse

OpenStudy (anonymous):

2x + y = 4 -2 -2

OpenStudy (ajprincess):

u have to subtract 2x both sides. 2x+y-2x=4-2x y=-2x+4 Nw compare this equation with y=mx+c and pls tell me the value of m. here m is the slope of the given line

OpenStudy (anonymous):

i did subtract 2x

OpenStudy (ajprincess):

after subtracting what did u get? @DvonCruse

OpenStudy (anonymous):

2

OpenStudy (ajprincess):

what is that 2?

OpenStudy (anonymous):

-2

OpenStudy (ajprincess):

is that the value of m?

OpenStudy (anonymous):

yea

OpenStudy (ajprincess):

ya it is right:)

OpenStudy (ajprincess):

let the required slope be m_1. the product of the slopes of the perpendicular lines is equal to -1. that is m*m_1=-1 -2*m_1=-1 so vat is m_1?

OpenStudy (anonymous):

thank you!

OpenStudy (ajprincess):

Welcome:) have u found m_1?

OpenStudy (ajprincess):

@DvonCruse

OpenStudy (anonymous):

m_1?

OpenStudy (ajprincess):

m_1 is the required slope